OpenStudy (anonymous):

Rewrite 2x^2+2y^2-8x+10y+2=0 in standard form. Find the center and radius of the circle.

OpenStudy (welshfella):

fist take 2 out = 2(x^2 + y^2 - 4x + 5y + 1) = 0 x^2 - 4x + y^2 + 5y + 1 = 0 we can now complete the square on the terms in x and in y can you do that ?

OpenStudy (anonymous):

do I combine like terms? @welshfella sorry if that's wrong im really confused

OpenStudy (welshfella):

you use the identity a^2 + 2ab = (a + b)^2 - b^2